Amy,Joe and David have a total of $135 in their wallets. Joe has $10 more than Amy. David has 3 times what Amy has. How much does each have ?
Answer by lenny460(1073)

Amy,Joe and David have a total of $135 in their wallets. Joe has $10 more than Amy. David has 3 times what Amy has. How much does each have ?

Let their wallets have:
Amy = x
Joe = x + 10 and
David = 3x

Therefore:

x + x + 10 + 3x = $135
2x + 3x + 10 = 135
5x + 10 = 135
5x = 125
x = $25

Answer:

Amy has $25, Joe has $35 and David has $75




Lennox Obuong
Algebra Tutor
Dallas, Texas